OpenStudy (anonymous):

\[\frac{ 3p ^{5} m ^{-2}}{ 18 PM ^{5} }\] simplify

OpenStudy (aum):

\[\Large \frac{ 3p ^{5} m ^{-2}}{ 18 pm ^{5} } = \frac{3}{18} * \frac{p^5}{p} * \frac{m^{-2}}{m^5} = \frac 16 * p^{5-1} * m^{-2-5} \] Simplify the exponents. Express result with positive exponents.

OpenStudy (anonymous):

i dont know how to express the result with positive exponents

OpenStudy (kropot72):

\[\frac{p ^{5-1}}{6m ^{5+2}}\]

OpenStudy (aum):

\[ \Large \frac 16 * p^{5-1} * m^{-2-5} = \frac 16 * p^{4} * m^{-7} = \frac 16\frac{p^{4}}{ m^{7} } \]